OpenStudy (anonymous):

derivative of sqrt(r^4 + 1)

OpenStudy (anonymous):

\[\sqrt(r^4 + 1) \]

OpenStudy (mertsj):

\[\frac{1}{2\sqrt{r^4+1}}(4r^3)\]

OpenStudy (anonymous):

and could you help show steps

OpenStudy (mertsj):